OpenStudy (anonymous):

Simplify. 1/i plz help

OpenStudy (anonymous):

Rationalize the denominator.

OpenStudy (anonymous):

\[\frac{1}{i} = \frac{1}{i}\ \cdot \frac ii = \frac{i}{i^2} = \frac{i}{-1} = \boxed{-i}\]This is something you need to memorize.

OpenStudy (anonymous):

Thank you